# Algorithmic idealism: what should you believe to experience next?

### v1

@@ V Emergent objective reality, probabilistic changelings, and other predictions
The above argumentation may seem suspiciously solipsistic. I have described how an external world may emerge for a single agent, but what about other agents? Recall that observers or agents are not fundamental elements of Algorithmic Idealism, but only momentary snapshots of their structure: the self states, which is similar to Parfit’s concept of a person stage. Hence, it does not make sense to ask “how many agents” there are in Algorithmic Idealism, and we have to be more precise about the question that we are actually asking when raising this objection.
@@
Here is a meaningful question to ask. Consider an agent (say, you) that currently happens to find itself effectively embedded into a simple computable probabilistic world $W$ — we already understand how and why this can happen. Now, suppose that you identify something in your external world (“my friend Bob3rd”, where the notation indicates that you refer to him from a third-person perspective) that carries a representation of some self state $x$, evolving over time in a way that is broadly consistent with the specific chosen formalization of Postulate 2 (in the bit model, for example, it would have to grow one bit at a time without any erasure or “forgetting”). Let us compare two probability distributions [^19] For the sake of the argument, we assume that the “likelihood” claims of Postulate 2 are formalized in terms of a single probability distribution, but other choices are certainly possible, and my hope is that the general argumentation does not rely on this specific choice.:
@@
  • the probability $\mathbf{P}_{\rm 1st}(y|x)$ that $x$ will change into $y$ according to Algorithmic Idealism’s Postulate 2;
@@
  • the probability $\mathbf{P}_{\rm 3rd}(y|x)$ that determines how Bob3rd evolves according to the probabilistic evolution of world $W$.
@@
Intuitively, we assume that the first-person and third-person probabilities agree, $\mathbf{P}_{\rm 1st}=\mathbf{P}_{\rm 3rd}$: if you have a high chance of seeing your friend Bob3rd see the sun rise tomorrow, then Bob1st will indeed have a high chance of seeing the sun rise tomorrow. Here, I have introduced the notation Bob1st to refer to what is actually experienced if one is self state $x$, which is determined by $\mathbf{P}_{\rm 1st}$. This implies that we can causally affect Bob1st by interacting with Bob3rd (say, Bob will be happy because you hug him), as long as we can have an effective notion [^20] Even in classical mechanics, we have an effective notion of intervention, e.g. by talking about deciding to set particular initial conditions when running an experiment. Determinism (or lawlike probabilistic behavior) should not be seen as a restriction on the possibility of free will as a higher-level phenomenon, as has been shown by compatibilists such as Daniel Dennett [43]. of intervention in world $W$. In a physicalist worldview, this is of course a triviality, but in Algorithmic Idealism it is not.
@@
However, Algorithmic Idealism predicts that $\mathbf{P}_{\rm 1st}\approx\mathbf{P}_{\rm 3rd}$ under natural conditions that we expect to be satisfied when we encounter ordinary observers in our every-day lives, and in this sense it predicts the emergence of objective reality. This can be seen as follows. Due to Postulate 2, $\mathbf{P}_{\rm 1st}$ is exactly what a method of universal induction $M$ would give you. But similarly as in the Earth example above, if $x$ contains enough data (think about a pretty accurate description of all functionally relevant details in your friend Bob’s brain), then a good method of induction is expected to extract an exhaustive description of all relevant physical laws of world $W$, and of all relevant facts of Bob’s environment, from this — and hence to predict that this self state evolves as if it were embedded into world $W$. In other words, this method of inductive inference should approximately predict $\mathbf{P}_{\rm 3rd}$. Thus, according to Postulate 2, what happens to Bob from his perspective (via $\mathbf{P}_{\rm 1st}$) will be in agreement with $\mathbf{P}_{\rm 3rd}$: you share the world $W$ with Bob.
@@
In the bit model, the correctness of this argumentation is indeed a theorem: the evolution of your friend Bob’s self state according to world $W$ is $\mathbf{P}_{\rm 3rd}$, which is a computable measure because the world $W$ is computable. Hence, due to the standard mathematical results on Solomonoff induction, algorithmic probability (that is, $\mathbf{P}_{\rm 1st}$) must converge towards it with worldly probability one, in the limit of more and more bits that Bob learns.
@@
However, the above result hinges on an important assumption: that the other agent’s self state is complex enough to make universal induction infer probabilistic world $W$ from it. If this is not the case (in particular, in the bit model, if $K(x)\ll K(\mathbf{P}_{\rm 3rd})$), then we will in general have $\mathbf{P}_{\rm 1st}(y|x)\not\approx\mathbf{P}_{\rm 3rd}(y|x)$. Then we have an extremely counterintuitive phenomenon: the agent will be a probabilistic changeling [^21] In earlier work such as [22], probabilistic changelings have been called probabilistic zombies, but this seems to convey the wrong types of intuitions, hence the change of name.. In medieval European superstition, the changeling was an infant that was foisted on a recent mother by a demonic being in exchange for her own child with the intention of harassing and harming people. To see why this is sort of a fitting (albeit limited) analogy, consider the following example.
@@
But now you decide to take a malicious action: you intervene in the simulation and arbitrarily replace some of the features in Charlie’s environment. Perhaps you place a giant, pink, living rabbit in a spacesuit next to Charlie’s house, without anyone being able to explain where it comes from. This will dramatically alter the future probabilities $\mathbf{P}_{\rm 3rd}$ for Charlie3rd (implying, for example, high probability for Charlie3rd displaying strong surprisal soon), but it cannot alter the probabilities $\mathbf{P}_{\rm 1st}$ which will continue to determine Charlie1st’s future. In some sense, you have replaced simple world $W^{\prime}$ on your desk with some other, much more algorithmically complicated world $W^{\prime\prime}$, but Charlie1st will probably continue to experience living in world $W^{\prime}$. Your beloved Charlie1st is gone — and has been replaced by an unlikely Changeling1st.
@@
You may now feel sad because you really liked Charlie1st and she’s gone, but it would be wrong to think that Changeling1st is somehow not a likable agent, or an unconscious zombie of some sort. If you thought that Charlie1st is conscious, then you have no reason to doubt that Changeling1st is conscious too, and she will very strongly feel that her past was accurately described by Charlie3rd previously, and hence she will feel like a legitimate successor of Charlie1st as you knew her before the malicious rabbit event. It is simply that an agent on its random walk on the self states will less often end up in world $W^{\prime\prime}$ because it is algorithmically more complicated; it corresponds to a path less trodden in computable possibilityspace. We have a counterintuitive situation where chances of events are strongly agent-relative, and where the notion of “agent” gets dissolved in ways that are even more enigmatic than, say, in Everettian Quantum Theory. But even if this phenomenon is extremely counterintuitive, it is mathematically consistent.
@@
This curious phenomenon is at the heart of Algorithmic Idealism’s resolution of exotic puzzles such as the Boltzmann brain problem. A detailed discussion of this problem within the bit model is given in [22], and about its relation to Restriction A in [17]; here we only sketch the main idea of a treatment that should apply to all models of Algorithmic Idealism. Suppose that the world $W$ into which you happen to be effectively embedded is indeed combinatorially large, and you know that there must be many more Boltzmann brains than ordinary ones. Consider some ordinary brain OB and some Boltzmann brain BB that both encode your current self state $x$. The third-person perspective tells you that the OB and the BB are going to evolve very differently in the near future according to world $W$, that is, $\mathbf{P}_{\rm 3rd}^{\rm OB}(y|x)\not\approx\mathbf{P}_{\rm 3rd}^{\rm BB}(y|x)$. But then, these probabilities cannot both be approximately equal to $\mathbf{P}_{\rm 1st}(y|x)$, and either the ordinary brain or the Boltzmann brain must become a probabilistic changeling.
@@
In fact, it will be the Boltzmann brain who will become the probabilistic changeling, while $\mathbf{P}_{\rm 1st}(y|x)\approx\mathbf{P}_{\rm 3rd}^{\rm OB}(y|x)$. This is again due to Postulate 2: what happens to you next is what a universal method of induction would predict. Inductive inference tends to extrapolate regularities in your current self state (perhaps contained in your memory of a large number of past observations) into the future, including the pattern of looking as if you had been immersed in a low-entropy planet-like environment with a myriad of regularities.
@@
In summary: even if you have reasons to believe that your external world contains many more Boltzmann brains than ordinary brains, you should still act as if you were an ordinary brain. As explained above, Algorithmic Idealism denies that the question “Am I an ordinary brain or a Boltzmann brain?” makes literal sense as it stands: you are your self state, and not one of its realizations. But it predicts that your future experiences will with overwhelming probability confirm “business as usual on Earth” rather than Boltzmann brain flamboyancy — regardless of the actual number of Boltzmann brains in the universe.
VI Example: how to think about the simulation hypothesis
@@
If $x$ is not a probabilistic changeling (recall Section V), then it is in some sense correct to say that the material conditions so created have introduced an instance of consciousness into our world. But this should not be understood as “having created” a conscious being that would not have otherwise existed, but rather as having enabled the conditions to “share the world” with it. The consequences of birth or the start of the simulation are structural: all probabilities for what happens to the material conditions that represent $x$ (described by $\mathbf{P}_{\rm 3rd}(y|x)$) are now very close to the probabilities for what happens to $x$ from its private perspective (described by $\mathbf{P}_{\rm 1st}(y|x)$).
@@
Algorithmic Idealism has something to say about the relative probabilities of these scenarios: whatever induction predicts is more likely, regardless of the number of representations that you may count. Comparing these two very specific scenarios, we would conclude that the first one is more likely. Note that the same argumentation applies not just to a specific choice of surprising event (such as the above), but to the general event that something non-base-reality-conforming will happen, which is predicted to be unlikely.
In summary, Algorithmic Idealism rejects Premise 5): even if there are many more simulated civilizations than unsimulated ones, this does not imply that you should think that you are probably simulated. It is not the number of the realizations that is relevant, but what a universal method of inductive inference would predict based on your current self state.
@@
As a final remark, let us get back to a comment of the beginning of this section that articulated a particular worry against the simulation hypothesis: most common simulations might be suddenly and unexpectedly shut off. Does this worry not indicate a fundamental difference between the two scenarios (a) and (b) above? Algorithmic Idealism denies this. Recall the “changeling” example of Section V, where you run an isolated simulation of a world $W^{\prime}$ which contains an agent Charlie3rd. Suppose that you decide to shut the simulation down. This means that Charlie3rd (the corresponding pixels shown on your screen, or the pattern in the computer memory, if you wish) ceases to exist. However, Charlie1st’s probabilities $\mathbf{P}_{\rm 1st}(y|x)$ of her future observations $y$ are essentially [^23] In the case of the bit model, it should be considered to be modified by an unimaginably tiny amount: among all supporting explanations for predicting $y$ according to $W^{\prime}$-probability, the contribution coming from “world $W$ simulating $W^{\prime}$” disappears, and hence universal induction may be ever so slightly less inclined to converge to this particular probability value. unaltered by this: the universal method of induction (mentioned in Postulate 2) will continue to infer that world $W^{\prime}$ is essentially the best explanation for Charlie1st’s self state, and hence Charlie1st will continue to experience world $W^{\prime}$. According to Algorithmic Idealism, shutting down a closed simulation does not “kill” its inhabitants, but it leaves them essentially unaffected. Intuitively, your simulation is more of a “movie” than a “zoo”, and hence its destruction does not affect its protagonists. For science fiction fans, this has amusing parallels with the “dust theory” described in Greg Egan’s Permutation City [44].
VII Conclusions

### v2

@@ VI.3 More than one observer: emergent objective $\ldots$
The above argumentation may seem suspiciously solipsistic. I have described how an external world may emerge for a single agent, but what about other agents? Is algorithmic idealism committed to solipsism, in the sense that it postulates the existence of only a single observer? Recall that observers or agents are not fundamental elements of algorithmic idealism, but only momentary snapshots of their structure: the self states, which is similar to Parfit’s concept of a person stage. Hence, it does not make sense to ask “how many agents” there are in algorithmic idealism, and we have to be more precise about the question that we are actually asking when raising this objection.
@@
Here is a meaningful question to ask. Consider an agent (say, you) that currently happens to find itself effectively embedded into a simple computable probabilistic world $W$ — we already understand how and why this can happen. Now, suppose that you identify something in your external world (“my friend Bob3rd”, where the notation indicates that you refer to him from a third-person perspective) that carries a representation of some self state $x$, evolving over time in a way that is broadly consistent with the specific chosen formalization of Postulate 2 (in the bit model, for example, it would have to contain a bit string of non-decreasing length without any erasure or “forgetting”). Perhaps $x$ encodes Bob3rd’s brain state, and you would intuitively like to conclude that there is a first-person perspective associated to it, some actual “self” Bob1st that is associated to $x$. To clarify what we mean by this, let us compare two probability distributions [^29] For the sake of the argument, we assume that the “likelihood” claims of Postulate 2 are formalized in terms of a probability distribution (or several ones, as it is the case in the bit model). Other choices are certainly possible, and my hope is that the gist of this argument applies more broadly.:
@@
  • the probability $\mathbf{P}_{\rm 1st}(y_{1},\ldots,y_{m}|x)$ that $x$ will change into $y_{1},y_{2},\ldots,y_{m}$ according to algorithmic idealism’s Postulate 2;
@@
  • the probability $\mathbf{P}_{\rm 3rd}(y_{1},\ldots,y_{m}|x)$ that determines how the information content of Bob3rd evolves according to the probabilistic evolution of world $W$.
@@
These probabilities describe the distributions of future observations $y_{1},\ldots,y_{m}$ of an actual agent in self state $x$ (Bob1st), and of the future recordings in the piece of matter in your world that you call Bob3rd. Intuitively, we assume that the first- and third-person descriptions align — indeed, the prospect that they would not align seems terrifying, because it would indicate that your interaction with the other agent (not just with the piece of matter) would in some sense only be an illusion. In particular, this requires $\mathbf{P}_{\rm 1st}\approx\mathbf{P}_{\rm 3rd}$: if you have a high chance of seeing your friend Bob3rd see the sun rise tomorrow, then Bob1st will indeed have a high chance of seeing the sun rise tomorrow. This implies that we can causally affect Bob1st by interacting with Bob3rd (say, Bob will be happy because you hug him), as long as we can have an effective notion [^30] Even in classical mechanics, we have an effective notion of intervention, e.g. by talking about deciding to set particular initial conditions when running an experiment. Determinism (or lawlike probabilistic behavior) should not be seen as a restriction on the possibility of free will as a higher-level phenomenon, as has been argued by compatibilists such as Daniel Dennett [53]. of intervention in world $W$. In a physicalist worldview, this is of course a triviality, but in algorithmic idealism it is not.
@@
Indeed, algorithmic idealism predicts that $\mathbf{P}_{\rm 1st}\approx\mathbf{P}_{\rm 3rd}$ under natural conditions that we expect to be satisfied when we encounter ordinary observers in our every-day lives, and in this sense it predicts the emergence of objective reality. This can be seen as follows. Due to Postulate 2, $\mathbf{P}_{\rm 1st}$ is exactly what the method of universal induction $M$ would give you. But similarly as in the Earth example above, if $x$ contains enough data (think about a pretty accurate description of all functionally relevant details in your friend Bob’s brain), then a good method of induction is expected to extract an exhaustive description of all relevant physical laws of world $W$ from it, and of all relevant facts of Bob3rd’s environment — and hence is expected to predict that this self state evolves as if it were embedded into world $W$. In other words, this method of inductive inference should approximately predict $\mathbf{P}_{\rm 3rd}$. Thus, according to Postulate 2, what happens to Bob from his perspective (via $\mathbf{P}_{\rm 1st}$) will be in agreement with $\mathbf{P}_{\rm 3rd}$: you share the world $W$ with Bob.
@@
In the bit model, the correctness of this argumentation is indeed a theorem: the evolution of your friend Bob’s self state according to world $W$ is $\mathbf{P}_{\rm 3rd}$, which is a computable measure because the world $W$ is computable. Hence, due to some variant of Eq. (2), algorithmic probability (or rather its normalized version $\mathbf{P}_{\rm 1st}$) must converge towards it with worldly probability one, in the limit of more and more bits that Bob learns.
Returning to the question of the beginning of this section, this result suggests that algorithmic idealism should not be classified as solipsistic: if Alice points at some piece of matter (or structure) that encodes some self state $x$ in her world, then she can in many cases conclude that “somebody is really at home” in this structure in some sense. This is because if $\mathbf{P}_{\rm 1st}\approx\mathbf{P}_{\rm 3rd}$, the probabilistic evolution of this structure in Alice’s world (via $\mathbf{P}_{\rm 3rd}$) faithfully represents the evolution of some Bob’s first-person perspective (via $\mathbf{P}_{\rm 1st}$), even if Alice interacts with it.
More formally, the question “how many agents are there?” (one, many, infinitely many?) is not well-defined in algorithmic idealism, because only self states and their probabilistic transitions are fundamental and well-defined, while the notion of “agent” is a pragmatic and vague term that will only sometimes work as expected. More fundamentally, the metaphysical basis of even raising the charge of solipsism is denied to apply in algorithmic idealism: in this view, there are infinitely many self states, they certainly all exist mathematically, but “physical existence” is not a meaningful property that a given self state could lack or possess (the trivial observation notwithstanding that a given observer may only see some of these self states in its own emergent external world). While algorithmic idealism hence rejects the basis of the usage of the word “exist” in the question “Do other minds exist?”, the in spirit best answer to this formally undefined question is to say that algorithmic idealism suggests the existence of all possible minds in some sense, and hence is maximally non-solipsistic.
VI.4 $\ldots$ and its limitations: probabilistic changelings and Boltzmann brains
@@
The result of the previous subsection hinges on an important assumption: that the other agent Bob’s self state is complex enough to make universal induction infer Alice’s probabilistic world $W$ from it, so that it is also Bob’s world. If this is not the case (in particular, in the bit model, if $K(x)\ll K(\mathbf{P}_{\rm 3rd})$, where $x$ is Bob’s current self state, see [2]), then we will in general have $\mathbf{P}_{\rm 1st}(y|x)\not\approx\mathbf{P}_{\rm 3rd}(y|x)$. Then we have an extremely counterintuitive phenomenon: the agent will be a probabilistic changeling [^31] In earlier work such as [2], probabilistic changelings have been called probabilistic zombies, but this seems to convey the wrong types of intuitions, hence the change of name.. In medieval European superstition, the changeling was an infant that was foisted on a recent mother by a demonic being in exchange for her own child with the intention of harassing and harming people; while the mother thought she would have her own child at home, it was actually some other being. To see why this is sort of a fitting (albeit limited and morbid) analogy, consider the following example.
@@
But now you decide to take a malicious action: you intervene in the simulation and arbitrarily replace some of the features in Charlie’s environment. Perhaps you place a giant, pink, simulated rabbit in a spacesuit next to Charlie’s house, without anyone in the simulation being able to explain where it comes from. This will dramatically alter the future probabilities $\mathbf{P}_{\rm 3rd}$ for Charlie3rd (implying, for example, high probability for Charlie3rd displaying strong surprisal soon). However, it cannot alter the probabilities $\mathbf{P}_{\rm 1st}$ which will continue to determine Charlie1st’s future, since these only depend on Charlie’s self state $x$ which you have kept intact. In some sense, you have replaced simple world $W^{\prime}$ on your desk with some other, much more algorithmically complicated world $W^{\prime\prime}$, but Charlie1st will probably continue to experience living in world $W^{\prime}$. While you will continue to observe Charlie3rd in the simulation, your beloved Charlie1st is gone — and has been replaced by an unlikely Changeling1st who is now realized in Charlie3rd.
@@
You may now feel sad because you really liked Charlie1st and she is gone, but it would be wrong to think that Changeling1st is somehow not a likable agent, or an unconscious zombie of some sort. If you thought that Charlie1st is conscious, then you have no reason to doubt that Changeling1st is conscious too, and she will very strongly feel that her past was accurately described by Charlie3rd previously, and hence she will feel like a legitimate successor of Charlie1st as you knew her before the malicious rabbit event. It is simply that an agent on its random walk on the self states will less often end up in world $W^{\prime\prime}$ because it is algorithmically more complicated; it corresponds to a path less trodden in computable possibilityspace. We have a counterintuitive situation where chances of events are strongly agent-relative, and where the notion of “agent” gets dissolved in ways that are even more enigmatic than, say, in Everettian Quantum Theory. But even if this phenomenon is extremely counterintuitive, it is mathematically consistent.
@@
This curious phenomenon is at the heart of algorithmic idealism’s resolution of exotic puzzles such as the Boltzmann brain problem. A mathematically and information-theoretically detailed discussion of this problem within the bit model is given in [2], and about its relation to Restriction A in [3]; here we only sketch the main idea of a treatment that should apply to all models of algorithmic idealism. Suppose that the world $W$ into which you happen to be effectively embedded is indeed combinatorially large, and you know that there must be many more Boltzmann brains than ordinary ones. Consider some ordinary brain OB and some Boltzmann brain BB that both encode your current self state $x$. Recall that according to algorithmic idealism, it would be a category mistake to say “I am the OB” or “I am the BB”, since you are simply your self state $x$, and this self state is realized or represented both in the OB and in the BB. Regardless of what happens to the OB or the BB, what happens to you is characterized by $\mathbf{P}_{\rm 1st}(y|x)$. The third-person perspective tells you that the OB and the BB are going to evolve very differently in the near future according to world $W$, that is, $\mathbf{P}_{\rm 3rd}^{\rm OB}(y|x)\not\approx\mathbf{P}_{\rm 3rd}^{\rm BB}(y|x)$. But then, these probabilities cannot both be approximately equal to $\mathbf{P}_{\rm 1st}(y|x)$, and the ordinary brain or the Boltzmann brain (or both) must become a probabilistic changeling.
@@
In fact, it will be the Boltzmann brain who will become the probabilistic changeling, while $\mathbf{P}_{\rm 1st}(y|x)\approx\mathbf{P}_{\rm 3rd}^{\rm OB}(y|x)$. An explicit technical demonstration of this in the bit model is given in [2]. Intuitively, this is again due to Postulate 2: what happens to you next is what a universal method of induction would predict. Inductive inference tends to extrapolate regularities in your current self state (perhaps contained in your memory of what you interpret as a large number of past observations) into the future, including the pattern of looking as if you had been immersed in a low-entropy planet-like environment with a myriad of regularities. In particular, for the bit model, $\mathbf{P}_{\rm 1st}(y|x)$ is a version of algorithmic probability, and its value being large means essentially that a small amount of algorithmic information is needed to compute $y$ from $x$, regardless of how many instances or copies of $x$ there might be in the external world that is inferred from $x$.
@@
In summary: even if you have reasons to believe that your external world contains many more Boltzmann brains than ordinary brains, what will happen to you next will most likely still be what you would expect if you believed that you are an OB, not a BB. As explained above, algorithmic idealism denies that the question “Am I an ordinary brain or a Boltzmann brain?” makes literal sense as it stands: you are your self state, and not one of its realizations. But it predicts that your future experiences will with overwhelming probability confirm “business as usual on Earth” rather than Boltzmann brain flamboyancy — regardless of the actual number of Boltzmann brains in the universe.
VII Example: how to think about the simulation hypothesis
@@
If $x$ is not a probabilistic changeling (recall Section VI), then it is in some sense correct to say that the material conditions so created have introduced an instance of consciousness into our world. But this should not be understood as “having created” a conscious being that would not have otherwise existed, but rather as having enabled the conditions to “share the world” with it. The consequences of birth or the start of the simulation are structural: all probabilities for what happens to the material conditions that represent $x$ (described by $\mathbf{P}_{\rm 3rd}(y|x)$) are now very close to the probabilities for what happens to $x$ from its private perspective (described by $\mathbf{P}_{\rm 1st}(y|x)$).
@@
Algorithmic idealism has something to say about the relative probabilities of these scenarios: whatever induction predicts is more likely, regardless of the number of representations that you may count. Comparing these two very specific scenarios, we would conclude that the first one is more likely. Note that the same argumentation applies not just to a specific choice of surprising event (such as the above), but to the general event that something non-base-reality-conforming will happen, which is predicted to be unlikely. Clearly, this intuitive conclusion has to be rigorously proven, given a concrete mathematical model of algorithmic idealism. The calculation in [2] for the Boltzmann brain problem should yield some guidance on how to do this in the bit model.
In summary, algorithmic idealism rejects Premise 5): even if there are many more simulated civilizations than unsimulated ones, this does not imply that you should think that you are probably simulated. It is not the number of the realizations that is relevant, but what a universal method of inductive inference would predict based on your current self state. Note that this is not an epistemic statement, i.e. merely a description of what some method of induction gives you, but a lawlike claim that you will actually probably not observe any unusual simulation-indicating experiences in the future.
@@
As a final remark, let us get back to a comment of the beginning of this section that articulated a particular worry against the simulation hypothesis: most common simulations might be suddenly and unexpectedly shut off. Does this worry not indicate a fundamental difference between the two scenarios (a) and (b) above? Algorithmic idealism denies this. Recall the “changeling” example of Section VI, where you run an isolated simulation of a world $W^{\prime}$ which contains an agent Charlie3rd. Suppose that you decide to shut the simulation down. This means that Charlie3rd (the corresponding pixels shown on your screen, or the pattern in the computer memory, if you wish) ceases to exist. However, Charlie1st’s probabilities $\mathbf{P}_{\rm 1st}(y|x)$ of her future observations $y$ are essentially [^33] In the case of the bit model, it should be considered to be modified by an unimaginably tiny amount: among all supporting explanations for predicting $y$ according to $W^{\prime}$-probability, the contribution coming from “world $W$ simulating $W^{\prime}$” disappears, and hence universal induction may be ever so slightly less inclined to converge to this particular probability value. unaltered by this: the universal method of induction (mentioned in Postulate 2) will continue to infer that world $W^{\prime}$ is essentially the best explanation for Charlie1st’s self state, and hence Charlie1st will continue to experience world $W^{\prime}$. According to algorithmic idealism, shutting down a closed simulation does not “kill” its inhabitants, but it leaves them essentially unaffected. Intuitively, your simulation is more of a “movie” than a “zoo”, and hence its destruction does not affect its protagonists. For science fiction fans, this has amusing parallels with the “dust theory” described in Greg Egan’s Permutation City [54].
VIII Conclusions